OpenStudy (anonymous):

Find the volume of the solid in R3 bounded by y=x^2, x=y2, z=x+y+21, and z=0.

OpenStudy (turingtest):

first step is t draw the region under the solid in the xy-plane, which is pretty easy in this case

OpenStudy (turingtest):

the region is bounded by\[0\le z\le x+y+21\]\[\sqrt x\le y\le x^2\]\[0\le x\le 1\]